May 2004

Sunday Monday Tuesday Wednesday Thursday Friday Saturday
 
 
 
 
 
 
1
DK Pipe pump-down test in progress.
2
DK Pipe pump-down test in progress.

Projected Work:
• Continue 5 Plane per day Installation, beginning at 6am daily.
• Continue Detector Commissioning, begining at 2:30pm daily.
• Begin installing rigging equipment for Absorber installation.

• Sat May 8 - elevator repair, expected to take 12 hrs

3
DK Pipe pump-down test completed - restore pipe to atm, slowly, through filter and dessicant to treat incoming air.

Five planes installed - 215, 214, 213, 212, 211. Plane 211 has cables attached and light-leak checked by 5pm.

Pass the Milestone for 25% of the Near Detector installed.

4
Five planes installed - 210, 209, 208, 207, 206. Plane 206 has cables attached and light-leak checked by 5pm.

Begin staging Contractor's tools in Absorber area.
5
Five planes installed - 205, 204, 203, 202, 201. Plane 201 has cables attached and light-leak checked by 5pm.

Continue staging Absorber installation equipment.
6
Five planes installed - 200, 199, 198, 197, 196. Plane 196 is on the detector by 11am.  Cables attached and light-leak checked by 4pm.

Continue staging Absorber installation equipment.  Also staging concrete shielding blocks to the hardstand.
7
Five planes installed - 195, 194, 193, 192, 191. Plane 191 is on the detector by 10:30am; cables attached and light-leak checked by 3pm.

Continue staging Absorber installation equipment.  Also staging concrete shielding blocks to the hardstand.
8
Repairs made to large elevator.
9
Projected Work:
• Continue 5 Plane per day Installation, beginning at 6am daily.
• Continue Detector Commissioning, begining at 2:30pm  or earlier daily.

Continue Absorber Installation.
10
Five planes installed - 190, 189, 188, 187, 186. Plane 186 is on the detector by 10am; Survey crew checking locations until 2pm; cables attached and light-leak checked by 5pm.

Test winch and cable to be used for Absorber installation.  Gantry crane set up.
11
Five planes installed - 185, 184, 183, 182, 181. Plane 181 is on the detector by 11am; cables attached and light-leak checked by 4pm.

First two concrete blocks set into place in Absorber Encl, located at the upstream end under the Decay Endcap.
12
Main elevator door gets stuck about 8am.  Repair contractor arrives about 10am and elevator is fixed by noon, but we were restricted to 6 people underground during this repair period.

Five planes installed - 180, 179, 178, 177, 176. Plane 176 is on the detector by 10:30am. Cabled and light-leak checked by 4:30.

Re-install Sump pump #2 at base of MINOS Shaft, between 1:30 and 4:40pm. Work performed by Absorber Inst crew.

Only Absorber work done today was to haul one C block up the hill.
13
Elevator door breaks again, restricting access until repair was completed at 10am.

Five planes installed - 175, 174, 173, 172, 171. Plane 171 is on the detector by 11am; cables attached and light-leak checked by 4pm.

C block set into place in Absorber Encl; 2 A blocks and another C block hauled up the hill, staged for tomorrow AM.
14
Five planes installed - 170, 169, 168, 167, 166. Plane 166 is on the detector by 11:00am. Cabled and light-leak checked by 5:30.

Continue installing Absorber concrete blocks.
15

 
 
 
 
16
Projected Work:
• Continue 5 Plane per day Installation, beginning at 6am daily.
• Continue Detector Commissioning, begining at 2:30pm or earlier daily.

Continue Absorner Installation
17
Five planes installed - 165, 164, 163, 162, 161. Plane 161 is on the detector by 11am; cables attached and light-leak checked by 4pm.

Electricians start work on grounding Detector readout racks.

Continue staging and installing Absorber concrete blocks. Lowered 2 As and 3 Bs.
18
Five planes installed - 160, 159, 158, 157, 156. Plane 156 is on the detector by 1:00pm. Delay due to difficulties mounting coil collar on a somewhat warped steel plate.  Cabled and light-leak checked by 5:30.

Electricians continue installing ground lines for Detector readout racks.

Continue installing Absorber concrete blocks.  Lowered 2 As and 3 Bs.  Also lowered the gantry cross-beam. Entire base level of blocks which surround DK pipe are in place.
19
Five planes installed - 155, 154, 153, 152, 151. Plane 151 is on the detector by 10am; cables attached and light-leak checked by 5pm.

At 10am while lowering plane 150 (next day's 1st plane) the shaft crane motor goes up in smoke; plane is left hanging in shaft.  At 2pm with assistance from crane repairman, the plane is brought back up and laid back on the floor of MSB.  

But the crane is not repaired.  

Electricians continue installing ground lines for Detector readout racks.

Tech begins installing connection hoses to FE racks for their water cooling.

Continue installing Absorber concrete blocks. Survey group measures as-founds on installed blocks and on DK Endcap.
20
No planes installed, and no Absorber work, due to shaft crane repair.

Parts required for crane repair are identified and ordered. Will arrive tomorrow.  Expect all repairs to be completed during Friday workday.

Electricians continue with miscellaneous jobs (in Detector Hall and elsewhere).

Detector commissioning work proceeded as usual.

Continue hose connections to readout racks.
21
Crane repaired by 12:30pm.

No planes installed.
Continue hose connections to readout racks.

No Absorber blocks delivered underground.
22

 
 
 
 
23
Projected Work:
• Continue 5 Plane per day Installation, beginning at 6am daily.
• Continue Detector Commissioning, begining at 2:30pm or earlier daily.

Continue Absorber Installation.
24
Five planes installed - 150, 149, 148, 147, 146. Plane 146 is on the detector by 10:30am.  Cabled and light-leak checked by 4:30.

Complete hose connections to readout racks.

Delivered 9 concrete blocks for Absorber. Concrete blocks on either side of DK pipe start being put in place.
25
Ten planes installed - 145, 144, 143, 142, 141. Plane 141 is on the detector by 9:30am; cables attached and light-leak checked by NOON. Then next 5 planes installed - 140, 139, 138, 137, 136. Plane 136 is on the detector by 4pm.  The N-V module on 136 is damaged when the cart lifting fixture is removed, after the plane is on the detector. Cables are NOT attached, as it is believed that the plane will have to be returned to New Muon for module replacement.

Delivered 5 concrete blocks and half of door which covers Hadron Monitor access slot.  Continue stacking mid-level of concrete blocks around DK pipe.
26
Techs decide that they can replace the N-V module in-place, in the Near Hall with the plane mounted on the detector stand. They fetch the spare N-V from Lab E, and perform the replacement, all by 8am. Cabled up and light-leak checked by 11:30am so today's planes can be installed.

Five planes installed - 135, 134, 133, 132, 131. Plane 131 is on the detector by 3pm.  Cabled and light-leak checked by 5:30.

Start leak-testing the Near Hall LCW water cooling system (used for racks & coil). Use pressurized air for initial leak checks.

Delivered 3 concrete blocks to Absorber Encl.  Completed stacking mid-level of concrete blocks around DK pipe. Blocks which form the Hadron Monitor access slot are in place - the door which covers the slot is not installed, as it is supported from other blocks not yet in place.
27
Five planes installed - 130, 129, 128, 127, 126. Plane 126 is on the detector by 10am.  Cabled and light-leak checked by 4:30.

Continue air-pressure leak test of Near Hall LCW system.

Delivered parts for 2nd gantry to Absorber Encl, to be used to unload Blu Blks at the top of the hill.  Start installing upper level of concrete blocks around DK pipe.
28
Last 5 planes of Spectrometer section due for installation today, but only get 3 done - planes 125, 124, 123, all bare steel.

Shaft crane motor makes smoke around 8am.  Repair called, but unavailable today - scheduled for tomorrow so as to have crane ready next Tuesday morning.

Continue air-pressure leak test of Near Hall LCW system. Start installing filters, etc for initial run of water in the system.

Continue installing top J blocks on Absorber shielding - these were already delivered to the Enclosure.
29

 
 
 
 
30

 
 
 
 
31
Memorial Day Holiday.